OpenStudy (anonymous):

(d/dx) of (2-t)(2^t)

OpenStudy (anonymous):

like this? \[\frac{d}{dx}[(2-t)(2^t)]\]

OpenStudy (anonymous):

wait...those are 2 different variables, are you sure thats the way the problem is posted? it should either be all 'x' or all 't'

OpenStudy (anonymous):

do yuo have the selected answers?

OpenStudy (anonymous):

usually y=f(x) and x=g(t) dy/dt= (dy/dt)(dx/dt)

OpenStudy (anonymous):

sorry it was all t and i got answer